Hi i need help with this question
solve 7+10log5(x)=13

Answer by rothauserc(4718)

7+10log5(x) = 13
subtract 7 from both sides of =
10log5(x) = 6
divide both sides of = by 10
log5(x) = 0.6
use definition of logarithm
x = 5^(0.6) = 2.626527804